OpenStudy (anonymous):

Will Medal and Fan How many solutions are there to the following system of equations? 2x + 6y = –12 10x + 32y = –62 0 2 1 infinitely many

OpenStudy (agl202):

First, lets reduce the equations.. 2x + 6y = -12 reduces to x + 3y = -6 10x + 32y = -62 reduces to 5x + 12y = - 31 so we have : x + 3y = -6 --->(-5) 5x + 12y = -31 ------------ -5x - 15y = 30 (result of multiplying by -5) 5x + 12y = -31 -------------add -3y = -1 y = 1/3 I am not going to figure x, because I do not need to. There is going to be 1 solution

OpenStudy (anonymous):

Wow... you are so smart :) Thank you so much, you are very kind.

OpenStudy (agl202):

Ur welcome... Glad I could help u! : )